OpenStudy (anonymous):

The slope of the line below is -3, and the y-intercept is (0, 2). What is the slope-intercept equation for the line?

OpenStudy (anonymous):

y = mx + b is slope intercept form. m = -3 b = 2

OpenStudy (anonymous):

medals to the people with the right answer

OpenStudy (anonymous):

A. y = -3x + 2 B. y = 2x - 3 C. y = 2x + 3 D. y = -3x - 2

OpenStudy (anonymous):

Kaiden, this is very easy. I just gave you all the info you need.

OpenStudy (anonymous):

y = mx + b m = -3 b = 2

OpenStudy (anonymous):

sub in those values for you equation.

OpenStudy (anonymous):

i dont understand this y = mx + b

OpenStudy (anonymous):

He just gave you the answer.. lol

OpenStudy (anonymous):

Plug in the number. y = mx + b m= -3 b= 2 y= -3x + 2

OpenStudy (anonymous):

Of course he gives you the medal because you gave him his answer. Derp.

OpenStudy (anonymous):

no he helped

OpenStudy (anonymous):

Even Klip would agree that I deserved the medal, you just gave it to him as payment for giving you the answer.

OpenStudy (anonymous):

LOL. Jcpd helped. He literally gave you the answer.

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

Anyway, let's not bag on him Klip. Just don't do that again, give medals to the people who deserve them.

OpenStudy (anonymous):

Thank you. :)